Jorgansen Lighting, Inc., manufactures heavy-duty street lighting systems for municipalities. The company uses variable costing for internal management reports and absorption costing for external reports to shareholders, creditors, and the government. The company has provided the following data:

Year 1 Year 2 Year 3
Inventories
Beginning (units) 200 170 190
Ending (units) 170 190 230
Variable costing net operating income $290,000 $279,000 $250,000

The company’s fixed manufacturing overhead per unit was constant at $570 for all three years.

Exercise 6-3 Part 1

Required:

1. Calculate each year’s absorption costing net operating income. (Enter any losses or deductions as a negative value.)

Answer is complete but not entirely correct.

Reconciliation of Variable Costing and Absorption Costing Net Operating Incomes
Year 1 Year 2 Year 3
Variable costing net operating income

$290,000

selected answer correct

$279,000

selected answer correct

$250,000

selected answer correct

Add (deduct) fixed manufacturing overhead deferred in (released from) inventory under absorption costing

17,100

selected answer incorrect

11,400

selected answer correct

22,800

selected answer correct

Absorption costing net operating income

$272,900

selected answer correct

$267,600

selected answer incorrect

$227,200

selected answer incorrect

2. Assume in Year 4 that the company’s variable costing net operating income was $250,000 and its absorption costing net operating income was $270,000.

a. Did inventories increase or decrease during Year 4?

  • Increase

  • Decrease

b. How much fixed manufacturing overhead cost was deferred or released from inventory during Year 4?

Fixed manufacturing overhead cost inventory during Year 4   
Answer #1
1
Year 1 Year 2 Year 3
Variable costing net operating income 290000 279000 250000
Add (deduct) fixed manufacturing overhead deferred in
(released from) inventory under absorption costing
(17100) 11400 22800
Absorption costing net operating income 272900 290400 272800
2a
As variable costing net operating income is less than absorption costing net operating income, inventories increase during the year
Increase
2b
Fixed manufacturing overhead cost deferred in inventory during Year 4 20000 (270000-250000)
Workings:
Year 1 Year 2 Year 3
Add (deduct) fixed manufacturing overhead deferred in
(released from) inventory under absorption costing
=(170-200)*570 =(190-170)*570 =(230-190)*570